### v1

# The Double-Ellipsoid Geometry of CLIP

Abstract
Contrastive Language-Image Pre-Training (CLIP) is highly instrumental in machine learning applications within a large variety of domains. We investigate the geometry of this embedding, which is still not well understood. We examine the raw unnormalized embedding and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin. We explain the benefits of having this structure, allowing to better embed instances according to their uncertainty during contrastive training. Frequent concepts in the dataset yield more false negatives, inducing greater uncertainty. A new notion of conformity is introduced, which measures the average cosine similarity of an instance to any other instance within a representative data set. We show this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector. Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
1 Introduction
Multi-modal approaches, particularly Contrastive Language-Image Pre-Training (CLIP) [44], have significantly advanced various computer vision tasks, enabling applications that were hard to imagine just a few years ago. These advancements include generating high-quality images from user prompts [45, 41]; performing open-vocabulary classification [18], segmentation [34, 57] and detection [54]; captioning [40, 7], and even semantic editing [29, 27]. Much of this success can be attributed to the key advancements made in multi-modal learning, with CLIP playing a key role. Following the remarkable success of vision-language models (VLMs), a wide range of adaptations has emerged for additional domains such as 3D [19, 5, 58], video [49, 37] and audio [53, 15].
Figure 1: Sketch of CLIP general geometry: image and text are embedded on linearly separable ellipsoid shells, not centered at the origin. This allows to control uncertainty in contrastive learning, where as themes become more rare (lower uncertainty) they reside farther from the mean modality vector.
Despite the practical advances, the structure of CLIP’s latent space is not fully understood. While properties such as alignment, uniformity, and the well-studied modality gap have been explored, they do not fully capture the geometry underlying this multi-modal space. A common approach is to investigate the geometry of the raw embedding after projection to the unit hypersphere. It was thus argued that CLIP features lie on a low aperture cone [35]. Since projection is an information reducing operation which “flattens” the data, we examine the raw embedding, before normalization to a unit norm. We show that CLIP’s raw latent space exhibits a double-ellipsoid geometry, one for text and one for images. Both ellipsoids are shifted from the origin (see Fig. 1). The offset from the origin explains the narrow cone effect, while their separability explains the modality gap phenomenon [35, 11, 47].
We perform a data-driven analysis of CLIP’s latent space using the MS COCO [36] validation set, without relying on any prior assumptions. We observe that both distributions of image and text follow the thin shell phenomenon [31, 32], typical of high dimensional data, in which most of the mass concentrates within some range from the mean.
We then examine the afforded advantages of this geometry. It is shown that shifting the mass from the origin allows to control the sharpness of the response in contrastive learning. A known phenomenon in unsupervised contrastive learning is the existence of false negatives [3, 33, 56]. That is, negative instances in the batch possessing similar concepts which should not be embedded far apart. False negatives can considerably degrade performance [4, 22]. This can be mitigated to some extent by reducing the sharpness of the response. Thus, as uncertainty grows the network produces a more “blurred” response, reducing the overall loss for the entire dataset. We refer to this as semantic blurring. Practically, this is done by embedding uncertain concepts closer (in terms of cosine distance) to the mean vector. Uncertainty grows as the concept is more frequent, with higher probability of false negatives in the batch. We thus predict that frequent concepts are better aligned to the mean vector of the ellipsoid. We validate this experimentally achieving excellent agreement with our hypothesis. Finally, we suggest a new way for image interpolation afforded by the deeper understanding of the geometry. Our contributions can be summarized as follows:
  1. 1.
    We investigate the raw CLIP embedding showing each modality resides on a separable ellipsoid shell not centered at the origin.
  2. 2.
    We analyze the advantages of such structure, allowing to control the sharpness of the response in contrastive learning.
  3. 3.
    False negative cases are shown to benefit mostly from this embedding. We demonstrate that the contrastive loss is optimized near CLIP’s ellipsoid offsets, for the MS-COCO dataset.
  4. 4.
    A new definition of concept conformity is proposed, quantifying the popularity of a concept. We find that conformity is correlated to (cosine) similarity to the mean vector with a remarkable Pearson correlation of 0.9998 (for both image and text on MS-COCO).
  5. 5.
    We reveal that the distribution of the conformity of image and text are different and are best matched in CLIP’s current positioning of the ellipsoids, leading to a possible explanation for the modality gap.
  6. 6.
    We demonstrate how conformity can be used to quantify and rank image and text generators.
  7. 7.
    Leveraging this geometry, we develop vertical SLERP (vSLERP), which extends interpolation capabilities in spontaneous, training-free image generation.
2 Related Work
Contrastive representation learning is a powerful learning scheme, where models are trained to associate positive pairs (e.g., different views of the same image [6]) closely in the embedding space while pushing negative pairs (e.g., different images) apart. This simple yet effective approach has led to significant advances across a wide range of applications, i.e. image classification [6, 17], natural language processing [14, 30], 3D analysis [1, 55] and more.
The latent space induced by contrastive learning has been widely explored [2, 24, 52, 51], often conceptualized as a normalized hypersphere [51, 35]. Alignment and uniformity [51] are key properties of the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss [6]. Optimizing alignment and uniformity was shown to be crucial for preserving rich semantic structures in the latent space, leading to improvements in downstream performance across multiple domains [11].
With the rise of cross-modal contrastive models, such as CLIP [44], which align images and text in a shared embedding space, new challenges in latent space geometry have emerged. A notable issue is the modality gap [35], where embeddings from different modalities, such as images and text, are separated in the shared latent space. Moreover, the narrow cone effect was observed [35, 47], where features occupy only a limited portion of the angular space.
One of the main challenges in multimodal contrastive learning is of obtaining high-quality pairs. Web-scale datasets may include mismatched positive pairs [9, 12, 38, 50] or mislabeled negative pairs that are actually positive, referred to as false negatives [3, 33, 56]. Numerous approaches have emerged to address this challenge, such as by identifying and introducing hard negative examples [4, 8, 46, 25]. Our observations are that false negatives appear to play a significant role in forming the geometry of CLIP’s latent space.
3 Random vectors in high dimensions
3.1 Notations
We investigate CLIP space induced by ViT-B/32 encoders of $n=512$ dimensions, $\mathcal{X}=R^{n}$. Let $\mathcal{X}_{i}\subset\mathcal{X}$ be the image subspace and $\mathcal{X}_{t}\subset\mathcal{X}$ be the text (captions) subspace. We will see that they are different and in fact linearly separable. Let $v\in\mathcal{X}$ be a vector in this space. We denote by $v_{i}\in\mathcal{X}_{i}$ vectors of images and by $v_{t}\in\mathcal{X}_{t}$ vectors of text. The symbol $\mathop{\operatorname{\mathbb{E}}}$ stands for the expected value. The respective modality mean of image and text are $m_{i}=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v_{i}\in\mathcal{X}_{i}\end{subarray}}[v_{i}]$ and $m_{t}=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v_{t}\in\mathcal{X}_{t}\end{subarray}}[v_{t}]$. Let $\tilde{v}$ be the vector after subtraction of the respective modality mean. That is, for images, $\tilde{v}_{i}=v_{i}-m_{i}\,:\,v_{i}\in\mathcal{X}_{i}$ and for text $\tilde{v}_{t}=v_{t}-m_{t}\,:\,v_{t}\in\mathcal{X}_{t}$.
Our statistical analysis and many experimental results are based on MS-COCO [36] validation set, a common standard image-text dataset.
3.2 High dimensional geometry of random vectors
It is often challenging to obtain good intuition on the probability manifold and its geometry in high dimensions. We outline below some fundamental concepts.
3.2.1 Thin shell theory
There is an intensive research related to the thin shell phenomenon [26, 43, 32, 23, 31]. Definitions of log concave distributions and isotropic random vectors appear in the supplementary. Since isotropic random vectors have a unit second moment for any $x(k),\,k=1,..n$, we get that the expected value of the squared Euclidean norm is
$\operatorname{\mathbb{E}}[\|x\|^{2}]=\operatorname{\mathbb{E}}[\sum_{k=1}^{n}x(k)^{2}]=\sum_{k=1}^{n}\operatorname{\mathbb{E}}[x(k)^{2}]=n.$ (1)
As shown for example in [43], $\operatorname{\mathbb{E}}[\|x\|^{2}]\approx\operatorname{\mathbb{E}}^{2}[\|x\|]$, the expected norm of $x$ can be approximated by
$\operatorname{\mathbb{E}}[\|x\|]\approx\sqrt{n}.$ (2)
For isotropic log-concave distributions we have the thin shell property:
Theorem 3.1 (Thin shell).
Let the thin shell parameter be defined by
$\sigma_{n}^{2}=\sup_{x}\mathop{\operatorname{\mathbb{E}}}(\|x\|-\sqrt{n})^{2},$
where the supremum is over isotropic, log-concave random vectors in $R^{n}$. Then $\sigma_{n}\leq c(\log n)^{\alpha}$, where $c$ is a universal constant.
Recent studies have shown this bound for $\alpha=4$ [32], $\alpha=2.23$ [23] and most recently for $\alpha=\frac{1}{2}$ [31]. See more details in the above papers and the references therein. Essentially, this means the mass of the distribution is concentrated around a shell of radius $\sqrt{n}$.
Let us farther examine this for the more general anisotropic case. Let $x=(x(1),...,\,x(n))$ be a vector of $n$ random variables of different distributions (not iid), each of mean zero. Let the norm of $x$, which is a random variable, be defined by $\|x\|=\mu_{norm}+y$, where $\mu_{norm}:=\operatorname{\mathbb{E}}[\|x\|]$ and $y$ is a random variable of zero mean. We examine the term $\operatorname{\mathbb{E}}[\|x\|^{2}]=tr(\mathbb{C})$, where $tr$ is the trace and $\mathbb{C}$ is the covariance matrix of $x$:
$\begin{array}[]{ll}\operatorname{\mathbb{E}}[\|x\|^{2}]&=\operatorname{\mathbb{E}}[(\mu_{norm}+y)^{2}]\\ =\operatorname{\mathbb{E}}[\mu_{norm}^{2}+2\mu_{norm}y+y^{2}]&=\mu_{norm}^{2}+\operatorname{var}(y).\end{array}$ (3)
Therefore, for $\mu_{norm}^{2}\gg\operatorname{var}(y)$ we can approximate
$\operatorname{\mathbb{E}}[\|x\|]=\mu_{norm}\approx\sqrt{\operatorname{\mathbb{E}}[\|x\|^{2}]}=\sqrt{tr(\mathbb{C})}.$ (4)
Here the squared expected Euclidean norm and the trace of the covariance matrix approximately coincide. We can thus view $\operatorname{std}(x(k))$ as a rescale of the coordinate system in dimension $k$, with respect to a unit sphere.
Figure 2: Normalized histograms of certain CLIP features. Image and text are clearly drawn from different statistics. On the right it is shown that even two features are sufficient to obtain full linear separability. The results of a linear SVM classifier are shown (blue dashed line, with $100\%$ accuracy on MS-COCO).
4 Geometric Analysis
We begin by examining the statistics of image and text in the CLIP embedding space $\mathcal{X}$. This part is completely data-driven without any prior assumptions related to the training process. We focus on the raw CLIP embedding, which is the output of the encoder before $L_{2}$ normalization, i.e. before projection onto the unit hypersphere. This projection loses important information. It basically “flattens” the original geometry artificially, in a manner which is hard to analyze. More details and statistical data are provided in the Supp.
Let us first examine the known modality gap phenomenon [35] in the raw embedding. In Fig. 2, normalized histograms are shown for features 93, 134 and 494 of the CLIP latent vector. We get a bimodal distribution where image and text are clearly not drawn from the same distribution. For feature 93, for instance, the KL-divergence between the distributions is $\approx 301$ (a value above 1 implies a considerable deviation between the distributions). It was previously shown in [48, 11] that image and text can be separated linearly. We find there are actually 9 features which serve as sort of “tags” for image and text. More formally, we can define the measure of separability of a feature $\ell$ by
$Sep(\ell)=\frac{|m_{i}(\ell)-m_{t}(\ell)|}{\sqrt{\operatorname{var}(v_{i}(\ell))+\operatorname{var}(v_{t}(\ell))}}.$ (5)
Figure 3: Separability of features (left) and 10 most significant features $\ell$ for image and text, with high absolute mean, compared to the feature’s standard deviation.
A plot of the features sorted by separability is shown in Fig. 3 (left). In Fig. 2 (right) it is shown that the modalities are linearly separable (with $100\%$ accuracy) using only two such tag features (93 and 134), based on a linear SVM classifier (decision boundary shown in blue). We can thus state the following property (which holds exactly at least for MS-COCO):
{mdframed}
[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 1: Image and text reside on separate subspaces, $\mathcal{X}_{i}\cap\mathcal{X}_{t}\approx\emptyset$.
Figure 4: Illustration of statistics of image and text features after subtracting the mean for each modality. Top, histograms of 10 first features of image (first row) and text (second row). Bottom, histograms of $\|\tilde{v}\|$ of images and of text. Although each feature separately has most of the volume near zero, the distribution of $\|\tilde{v}\|$ has no volume below some threshold, indicating a thin shell phenomenon, typical for high dimensions.
In Fig. 4, we show some statistics of the features of $\tilde{v}_{i}$ and $\tilde{v}_{t}$ (where the mean is subtracted). The first 10 features in each vector are shown for both modalities. The distribution appears smooth, unimodal, with peak around zero. The norm $\|\tilde{v}\|$, however, is distributed within a small range (thin shell) such that there is no mass near zero.
We can further check the validity of Eq. 3, we examine images here. In the case of MS-COCO statistics we have: $\mu_{norm}=7.5873$, $\operatorname{var}(y)=0.1914$, yielding $\mu_{norm}^{2}=57.5671\gg\operatorname{var}(y)$, where the approximation $\sqrt{\operatorname{\mathbb{E}}[\|x\|^{2}]}=7.6007$ is just with $0.18\%$ relative error. We can therefore conclude: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 2: The mass of each modality is concentrated within a thin shell, with zero mass near the mean of the distribution.
Figure 5: Normalized histograms of the variance of features (two left plots) indicates a long tail of high variance, hence an ellipsoid rather than an hypersphere. Off diagonal dominance, Eq. 6, implies certain strong feature correlations, or a tilted ellipsoid.
Let us now investigate the geometry of each shell. We examine the variance of each feature $\ell$. In a uniform hypersphere embedding we expect to have similar variance for all dimensions. We observe in Fig. 5 (left part) this is not the case, with a long tail distribution, where some features exhibit considerably larger variance, hence an ellipsoid structure: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 3: The embedding of both text and image is of an ellipsoid shell.
We now examine inter-correlations between features. Let us define off-diagonal dominance of a row $\ell$ in the covariance matrix $\mathbb{C}$ by
$ODD(\ell)=\frac{\sum_{k\neq\ell}|\mathbb{C}_{\ell k}|}{\mathbb{C}_{\ell\ell}}.$ (6)
Diagonally dominant matrices have $ODD(\ell)<1,\,\forall\ell$ ensuring a non-singular matrix. We observe (see Fig. 5 two right plots) that the off diagonals are significant, implying non-negligible correlation between features, thus: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 4: The ellipsoids of both modalities are tilted.
Finally, we check the location of each ellipsoid, with respect to the origin. We recall $m_{i}$, $m_{t}\in R^{n}$ are the mean value vectors of image and text. Let $\sigma_{i}$, $\sigma_{t}\in R^{n}$ be the standard deviation vectors of image and text, respectively. We have $\frac{\|m_{i}\|}{\|\sigma_{i}\|}=0.94$ and $\frac{\|m_{t}\|}{\|\sigma_{t}\|}=1.03$. Viewing $\|\sigma\|$ as a mean vector magnitude of the ellipsoid shell, the means are significantly shifted from the origin, compared to the size of the ellipsoid. This is caused by a few features, with strong deviation from the origin (compared to the respective feature’s standard deviation), as shown in Fig. 3 (middle and right). Thus we can state: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 5: The ellipsoids are not centered near the origin.
A linear transformation can be applied to the embedding to transform the ellipsoid into a sphere. This process is called whitening [10, 28]. Analysis of whitened CLIP is in the supplementary, providing indications that multivariate normal distribution is a rather crud approximation.
5 Loss behavior on a double-ellipsoid
In this section, we validate that a shifted double-ellipsoid structure achieves optimality in terms of the CLIP contrastive learning loss.
For a batch containing $M$ image-text pairs, we denote by $\bar{v}_{i}^{j}=\frac{v^{j}_{i}}{\|v^{j}_{i}\|}$ and $\bar{v}_{t}^{j}=\frac{v^{j}_{t}}{\|v^{j}_{t}\|}$ the normalized image and text features of the $j$-th pair in the batch respectively. The multi-modal learning loss used in CLIP is the normalized temperature-scaled cross entropy loss (NT-Xent), a variation of InfoNCE [42] loss:
$\ell_{clip}:=-\frac{1}{2}\mathop{\mathbb{E}}_{\begin{subarray}{c}j,k\in M\end{subarray}}\left[\log\frac{e^{{\bar{v}_{t}^{j\top}}\bar{v}^{j}_{i}/\tau}}{\sum_{j}e^{\bar{v}_{t}^{j\top}\bar{v}_{i}^{k}/\tau}}+\log\frac{e^{{\bar{v}_{t}^{j\top}}\bar{v}^{j}_{i}/\tau}}{\sum_{j}e^{\bar{v}_{t}^{k\top}\bar{v}_{i}^{j}/\tau}}\right].$ (7)
As observed by [51], the loss can be decomposed into two terms: (1) Alignment, which encourages high cosine similarity for positive pairs, and (2) Uniformity, which encourages low cosine similarity among negative samples.
$\begin{split}\ell_{clip}&:=-\overbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}j\in M\end{subarray}}[\bar{v}_{t}^{j\top}\bar{v}^{j}_{i}/\tau]}^{\text{alignment}}+\\ &\overbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}k\in M\end{subarray}}\left[\frac{1}{2}\log\sum_{j=1}^{M}e^{\bar{v}_{t}^{j\top}\bar{v}^{k}_{i}/\tau}+\frac{1}{2}\log\sum_{j=1}^{M}e^{\bar{v}_{t}^{k\top}\bar{v}_{i}^{j}/\tau}\right]}^{\text{uniformity}}.\end{split}$ (8)
To empirically analyze the uniformity and alignment terms in Eq. 8 alongside the overall loss in Eq. 7, we use the MS-COCO validation set. Fig. 6 shows the overall loss (bottom) and its breakdown into uniformity and alignment losses (top). We treat the entire validation set (5k samples) as a single batch. The overall loss is further separated into correctly classified, misclassified, and combined cases; the union of the correct and misclassified is equivalent to both, and they are mutually exclusive.
Figure 6: Loss with respect to position of center of embedding. The parameter $\alpha$ controls the center of embedding, see Eq. 9 ($\alpha=0$ is the current non-origin-centered position of CLIP). (Top). For non-origin-centered position, the unified loss is optimally balancing between uniformity and alignment. (Bottom). The loss is ascending when only misclassified instances are examined, and descending for well-classified. The overall accuracy, contains both, well balanced at $\alpha\approx 0$.
In this experiment, we examine different values of the mean value of the image embedding. For simplicity, we apply linear interpolation and extrapolation of the mean relative to the origin, using a single scalar parameter $\alpha$. This measure is conducted on a grid of $\alpha$ values from -1 to 1, with the loss calculated on image features as follows:
$v_{i}^{j^{\prime}}=v_{i}^{j}-\alpha\cdot m_{i}\quad\forall j\in M.$ (9)
The values of $v_{t}$ remain unchanged.
The results show that the loss for correctly classified samples decreases monotonically with the shift toward the origin (i.e. that for perfect alignment as assumed for example by [35], shifting to the origin would be preferable). Conversely, the loss for misclassified samples increases. The overall loss balances alignment and uniformity for both correctly and misclassified samples, reaching an optimal $\alpha$ near zero. This aligns with the current CLIP embedding, though some deviation is expected, as the MS-COCO validation set is only an approximation of the full training set.
For completeness, the supplementary material includes the same experiment with the text ellipsoid shifted instead of the image, showing consistent behavior. To conclude:
{mdframed}
[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 6: CLIP’s contrastive loss is optimized for shifted ellipsoids, balancing alignment and uniformity for both correctly classified and misclassified instances.
6 False Negatives and Conformity
We show here how the geometric structure of the embedding, found in previous sections, is advantageous when false negatives exist. We then propose the notion of conformity, which appears to play a major role in forming the distribution of the latent space.
In contrastive learning, a well known problem is the phenomenon of false negatives. Essentially, concepts which are not dedicated pairs have similar meaning. Thus, they do not represent well negative examples and should not be embedded far apart. This problem appears for both single and multi-modality cases. There have been several suggestions to solve this problem by introducing new learning procedures or proposing alternative contrastive losses [4, 8].
In the case of CLIP, however, the training is performed with a contrastive loss which does not take into consideration directly the false negative problem. We claim that the issue is mitigated, to some extent, by a specialized embedding. In classification and segmentation, as uncertainty grows, the network yields softer values, representing the lower membership probability to the class. In Fig. 7 (top) a typical segmentation score of semantic segmentation is shown, where the boundary of the sheep is blurred, due to reduced confidence. We refer to this as semantic blur. For a contrastive network, when false negatives are suspected we expect lower confidence and a more blurred response. In the case of a high dimensional sphere centered at the origin, blurring is not trivial. Small perturbations yield large disparities in cosine distance. We show below how a sphere not centered at the origin can mitigate this.
Figure 7: Top: blur of the segmentation score (right), typical in sematic segmentation, as the uncertainty of object-membership grows. Bottom: the similarity histograms of normally distributed samples are shown for the mean vector (blue) and a vector furthest from the mean (“extreme”, orange). The plots are done for a sphere centered near the origin (left) and for one with a center far from the origin (right). See full details of the experiment in the text. In the case of contrastive learning, blur can be achieved, as uncertainty grows, by having a sphere centered far from the origin. The desired blur can be controlled by positioning the embedded vector closer to the sphere center (blur) or farther from it (sharp).
[…]
{mdframed}
[ middlelinecolor = blue, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Prediction 1: Common themes, which occur more frequently in the training set, are expected to be embedded in closer proximity to the mean vector.
Figure 8: High and low conformity of natural images from MS-COCO [36]. Low-conformity images often depict unique, distinguishable individuals or objects, whereas high-conformity images capture common scenes that could be found anywhere.
6.1 Conformity
In order to validate our prediction, we first formalize the term common themes, by defining a new notion, termed conformity.
Definition 1 (Conformity).
Conformity of a vector $v^{j}$ within a set $S$ measures the expected value of the cosine similarity to this vector:
$C(v^{j})=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v^{k}\in S\\ j\neq k\end{subarray}}[\cos(v^{j},v^{k})],$ (10)
where for a given finite set $S$, the empirical mean is taken.
To provide more intuition, we present examples of high and low conformity from MS-COCO in Fig. 8, as well as on ImageNet-a [21] and ImageNet-R [20] in the Supp. Following our prediction above, we propose a surrogate measure of conformity (which is much faster to compute). The estimation uses the following definition.
Definition 2 (Estimated conformity).
In contrastive learning embedding, for a given set of vectors $S$ with mean $m=\mathop{\operatorname{\mathbb{E}}}_{\begin{subarray}{c}v\in S\end{subarray}}[v]$, the estimated conformity of $v^{j}\in S$ is:
$\hat{C}(v^{j})=a\cdot\cos(m,v^{j})+b,$ (11)
where $a$ and $b$ are scalars determined by the specific embedding.
Figure 9: Conformity. Estimated conformity $\hat{C}$, Eq. 11, against conformity $C$, Eq. 10, on MS-COCO [36]. The correlation is almost perfect. We can thus use the proposed estimated conformity reliably to quantify how common a sample is. Employing $\hat{C}$ is much faster, using a single mean vector $m$, without the need to store an entire set of representative image-text pairs. We also show a few caption examples of high and low conformity, with the same number of words (8). More exotic captions have lower conformity.
In Fig. 9, $C$ versus $\hat{C}$ are plotted for the entire MS-COCO set, for both image and text embeddings. A close to perfect correlation is obtained, with Pearson correlation of 0.9998 for both image and text where $a=1.411$, $b=-0.008$ for text and $a=1.461$, $b=-0.002$ for images.
6.2 Modality gap assists in distribution matching
We now aim to provide a reason that can justify the presence of the well known modality gap [35]. Our rationale for that phenomenon is as follows.
Figure 10: Conformity Differences. The conformity distributions of text and image modalities differ, as a common image may be described by a unique caption, and vice versa.
The same incentive of having a mean not centered at the origin applies for both image and text modalities. However, in a single image-pair instance the uncertainty for each modality may differ (see Fig. 10). The same arguments as before promote uncertain instances to be near the mean and certain ones to be far from it. If both image and text of a pair are embedded at the same location - we may get contradicting requirements. Having separate embeddings for text and image allows to control the uncertainty of each instance for each modality. More generally, we would like to match the distribution of the conformity of both modalities. In Fig. 11 we show the KL-divergence of the conformity distribution as a function of $\alpha$, a parameter controlling the distance of the mean from the origin, as in Eq. 9, see illustration in Fig. 6. We show that the best distribution match is near $\alpha=0$, that is, in the current embedding of CLIP.
Figure 11: Matching the distribution of conformity for image and text by introducing a modality gap. The parameter $\alpha$ controls the shift from the origin for images (as in Fig. 6). Both image and text conformity distributions are aligned when $\alpha\approx 0$, i.e. without any shift from the trained setting. $KL_{\alpha=0}\approx 0.14$ yields a good distribution matching.
Figure 12: Vertical SLERP (vSLERP) - Optimization-free, spontaneous semantic editing. Images interpolated with vSLERP preserve the same object with slight pose variations, while roughly maintaining backgrounds. The interpolation magnitude is controlled by $\alpha$.
Figure 13: Conformity analysis of Captioning and Image Synthesis. Image Synthesis (top): Glide [41] tends to generate more common images, often struggling with fine-grained details, whereas unCLIP [45] generates images with more delicate details, aligning more closely with the distribution of natural images. Though there is still room for improvement. Captioning (bottom): ClipCap [40] generates captions with higher conformity (i.e., more common descriptions) compared to human-annotated captions. In contrast, Caption Reward [7] produces captions with lower conformity, even surpassing human annotations in terms of uniqueness.
7 Applications
7.1 Conformity as a measure of expressiveness
We propose to use (estimated) conformity as a metric to assess the richness of generative methods. Specifically, we measure the conformity of images generated from MS-COCO captions by unCLIP [45] and Glide [41], visualizing the results in Fig. 13. Glide-generated images exhibit high conformity, indicating less detail and diversity, whereas unCLIP images are more varied and detailed. However, both models fall short of matching the diversity found in real images.
Similarly, we evaluate captioning methods by measuring conformity in captions generated using ClipCap [40] and Caption Reward [7]. ClipCap tends to produce common and trivial captions, while Caption Reward generates captions with diversity surpassing even human annotations.
7.2 Unguided, training-free semantic generation
The unCLIP framework [45] introduces an image interpolation technique using spherical linear interpolation (SLERP) to transform a source image into a target image gradually. While this method produces visually appealing results, it often fails to preserve the same instance throughout interpolation, instead generating random instances.
In Fig. 12, we show images generated by an extension of SLERP, which we term as vertical SLERP (vSLERP), defined as:
$\textit{vSLERP}(v^{j},v^{k},\theta_{0},\alpha)=\textit{SLERP}(v^{j}-\alpha m,v^{k}-\alpha m,\theta_{0})+\alpha m$ (12)
For brevity, $m_{i}$ and $v_{i}$ are referred to as $m$ and $v$. With a fixed $\Theta=\Theta_{0}$, adjusting $\alpha$ allows controlled manipulation of the same instance (values in the range $-2\leq\alpha\leq 2$ tend to produce visually appealing images). This approach parallels real-image editing techniques; however, unlike methods relying on text inversion [16, 13, 39] or test-time optimization [27], which are computationally heavy, vSLERP requires no training or optimization, thus, highly efficient.
8 Discussion and Conclusion
This paper analyzes the raw CLIP embeddings before projection onto a unit hypersphere, revealing that each modality’s distribution forms a distinct, shifted ellipsoid. These ellipsoids differ in properties such as their center of mass and radii and are disjoint, hence the modality gap phenomenon [35]. Upon projection onto a unit hypersphere, they occupy only a limited angular space, which appears as a narrow cone effect [35, 47]. We introduce the concept of conformity, a measure of similarity with the entire set, to gain further insights into the double-ellipsoid geometry. We observe that each modality has a unique conformity distribution, with optimal matching achieved when the ellipsoids are positioned away from the center. Using this framework, we evaluate the diversity of captioning and image synthesis methods and propose a training-free interpolation technique, vertical SLERP (vSLERP), for specific object interpolation.

### v2

@@
 Contrastive Language-Image Pre-Training (CLIP) is highly instrumental in machine learning applications within a large variety of domains.
-We investigate the geometry of this embedding, which is still not well understood.
-We examine the raw unnormalized embedding and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin.
+We investigate the geometry of this embedding, which is still not well understood, and show that text and image reside on linearly separable ellipsoid shells, not centered at the origin.
 We explain the benefits of having this structure, allowing to better embed instances according to their uncertainty during contrastive training.
@@
 Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
-1 Introduction
-Multi-modal approaches, particularly Contrastive Language-Image Pre-Training (CLIP) [44], have significantly advanced various computer vision tasks, enabling applications that were hard to imagine just a few years ago.
-These advancements include generating high-quality images from user prompts [45, 41]; performing open-vocabulary classification [18], segmentation [34, 57] and detection [54]; captioning [40, 7], and even semantic editing [29, 27].
-Much of this success can be attributed to the key advancements made in multi-modal learning, with CLIP playing a key role.
-Following the remarkable success of vision-language models (VLMs), a wide range of adaptations has emerged for additional domains such as 3D [19, 5, 58], video [49, 37] and audio [53, 15].
+Machine Learning, Vision Language Models, Modality Gap, ICML
 Figure 1: Sketch of CLIP general geometry: image and text are embedded on linearly separable ellipsoid shells, not centered at the origin.
 This allows to control uncertainty in contrastive learning, where as themes become more rare (lower uncertainty) they reside farther from the mean modality vector.
-Despite the practical advances, the structure of CLIP’s latent space is not fully understood.
-While properties such as alignment, uniformity, and the well-studied modality gap have been explored, they do not fully capture the geometry underlying this multi-modal space.
-A common approach is to investigate the geometry of the raw embedding after projection to the unit hypersphere.
-It was thus argued that CLIP features lie on a low aperture cone [35].
-Since projection is an information reducing operation which “flattens” the data, we examine the raw embedding, before normalization to a unit norm.
-We show that CLIP’s raw latent space exhibits a double-ellipsoid geometry, one for text and one for images.
-Both ellipsoids are shifted from the origin (see Fig.
-1).
-The offset from the origin explains the narrow cone effect, while their separability explains the modality gap phenomenon [35, 11, 47].
-We perform a data-driven analysis of CLIP’s latent space using the MS COCO [36] validation set, without relying on any prior assumptions.
-We observe that both distributions of image and text follow the thin shell phenomenon [31, 32], typical of high dimensional data, in which most of the mass concentrates within some range from the mean.
-We then examine the afforded advantages of this geometry.
-It is shown that shifting the mass from the origin allows to control the sharpness of the response in contrastive learning.
-A known phenomenon in unsupervised contrastive learning is the existence of false negatives [3, 33, 56].
-That is, negative instances in the batch possessing similar concepts which should not be embedded far apart.
-False negatives can considerably degrade performance [4, 22].
-This can be mitigated to some extent by reducing the sharpness of the response.
-Thus, as uncertainty grows the network produces a more “blurred” response, reducing the overall loss for the entire dataset.
-We refer to this as semantic blurring.
-Practically, this is done by embedding uncertain concepts closer (in terms of cosine distance) to the mean vector.
-Uncertainty grows as the concept is more frequent, with higher probability of false negatives in the batch.
-We thus predict that frequent concepts are better aligned to the mean vector of the ellipsoid.
-We validate this experimentally achieving excellent agreement with our hypothesis.
-Finally, we suggest a new way for image interpolation afforded by the deeper understanding of the geometry.
-Our contributions can be summarized as follows:
+1 Introduction
+Multi-modal approaches, particularly Contrastive Language-Image Pre-Training (CLIP) (Radford et al., 2021), have revolutionized computer vision tasks, enabling applications such as high-quality image generation (Ramesh et al., 2022; Nichol et al., 2021), open-vocabulary classification (He et al., 2023), segmentation (Liang et al., 2023; Yu et al., 2024), detection (Wu et al., 2023), captioning (Mokady et al., 2021; Cho et al., 2022), and semantic editing (Kim et al., 2022; Kawar et al., 2023).
+Beyond images, CLIP’s success extends to 3D (Hegde et al., 2023; Chen et al., 2023; Zhang et al., 2022), video (Tang et al., 2021; Luo et al., 2022), and audio domains (Wu et al., 2022; Guzhov et al., 2022).
+Despite these advances, the structure of CLIP’s latent space remains poorly understood.
+Existing studies focus on properties like alignment, uniformity, and the modality gap (Liang et al., 2022) but overlook the geometry underlying this multi-modal space.
+The L2-normalization phase, which is integral when performing cosine similarity, practically reducing the data to the unit hypersphere.
+Since normalization is an information-reducing process, understanding the primary embeddings prior to normalization can reveal deeper insights into the latent space geometry.
+In this paper, we propose analyzing the pre-normalized CLIP primary embedding for three key reasons: (1) Enhancing downstream tasks.
+While $L^{2}$-normalization is integral to the cosine similarity used during training, the primary embedding is directly employed in critical downstream tasks, including image generation and semantic editing.
+Analysis of the latent geometry can enhance the performance of these tasks.
+(2) Semantic significance of magnitude.
+Despite the cosine similarity is agnostic to the norm, we observe that magnitude still plays a significant and meaningful role.
+Notably, the largest embeddings in MS-COCO correspond to unusual or exotic captions (e.g., “I am not sure what this image is”, see full histogram and examples in Figure 14 in the Appendix).
+(3) Deeper understanding of contrastive learning.
+CLIP is an exceptional semantic encoder achieved through a rather generic contrastive loss and huge training data.
+Investigating the solutions found by CLIP allows deeper insights on contrastive learning, possible approaches to tackle false negatives and may shed light on unresolved phenomena, such as the modality gap and the narrow cone effect (Liang et al., 2022).
+Our analysis reveals that CLIP’s primary latent space exhibits a double-ellipsoid geometry, with one ellipsoid for images and another for text.
+Both are shifted from the origin (see Fig.
+1), in line with the narrow cone effect and the modality gap (Liang et al., 2022; Fahim et al., 2024; Schrodi et al., 2024).
+Using the MS-COCO validation set (Lin et al., 2014), we show that both modalities exhibit the thin-shell phenomenon (Klartag, 2023; Klartag & Lehec, 2022), where most of the mass concentrates within a specific range from the mean.
+This geometry affords several advantages.
+The offset from the origin allows CLIP to control the sharpness of its response in contrastive learning, mitigating false negatives (Byun et al., 2022; Li et al., 2022; Yang et al., 2022);instances that are conceptually similar but incorrectly treated as negatives.
+Frequent concepts with higher uncertainty are embedded closer to the mean vector, a phenomenon we term semantic blurring, reducing loss and improving performance.
+Our experiments confirm that frequent concepts are better aligned to the mean vector of the ellipsoid, achieving excellent agreement with our hypothesis.
+Leveraging this deeper understanding, we introduce a new definition of concept conformity, quantifying how close a sample resides with respect to all others.
+We find a remarkable correlation between conformity and cosine similarity to the mean vector (Pearson correlation: 0.9998 for MS-COCO).
+Furthermore, we show that the distribution of conformity differs between modalities, with CLIP’s ellipsoid alignment offering a plausible explanation for the modality gap.
+Our contributions are as follows:
   1. 1.
-    We investigate the raw CLIP embedding showing each modality resides on a separable ellipsoid shell not centered at the origin.
+    We reveal that CLIP embeddings form separable ellipsoid shells for each modality, shifted from the origin.
   2. 2.
-    We analyze the advantages of such structure, allowing to control the sharpness of the response in contrastive learning.
+    We analyze the benefits of this structure, including its role in controlling sharpness in contrastive learning.
   3. 3.
-    False negative cases are shown to benefit mostly from this embedding. We demonstrate that the contrastive loss is optimized near CLIP’s ellipsoid offsets, for the MS-COCO dataset.
+    We show that frequent concepts benefit most from this geometry, optimizing the contrastive loss near the ellipsoid offsets for MS-COCO.
   4. 4.
-    A new definition of concept conformity is proposed, quantifying the popularity of a concept.
-We find that conformity is correlated to (cosine) similarity to the mean vector with a remarkable Pearson correlation of 0.9998 (for both image and text on MS-COCO).
+    We define concept conformity and demonstrate its strong correlation with similarity to the mean vector, offering insights into semantic organization.
   5. 5.
-    We reveal that the distribution of the conformity of image and text are different and are best matched in CLIP’s current positioning of the ellipsoids, leading to a possible explanation for the modality gap.
+    We highlight the role of conformity in explaining the modality gap and propose its use in ranking text and image generators.
   6. 6.
-    We demonstrate how conformity can be used to quantify and rank image and text generators.
-  7. 7.
-    Leveraging this geometry, we develop vertical SLERP (vSLERP), which extends interpolation capabilities in spontaneous, training-free image generation.
+    We introduce vertical SLERP (vSLERP), an interpolation method leveraging the geometry of CLIP’s latent space.
 2 Related Work
-Contrastive representation learning is a powerful learning scheme, where models are trained to associate positive pairs (e.g., different views of the same image [6]) closely in the embedding space while pushing negative pairs (e.g., different images) apart.
+Contrastive representation learning is a powerful learning scheme, where models are trained to associate positive pairs (e.g., different views of the same image (Chen et al., 2020)) closely in the embedding space while pushing negative pairs (e.g., different images) apart.
 This simple yet effective approach has led to significant advances across a wide range of applications, i.e.
-image classification [6, 17], natural language processing [14, 30], 3D analysis [1, 55] and more.
-The latent space induced by contrastive learning has been widely explored [2, 24, 52, 51], often conceptualized as a normalized hypersphere [51, 35].
-Alignment and uniformity [51] are key properties of the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss [6].
-Optimizing alignment and uniformity was shown to be crucial for preserving rich semantic structures in the latent space, leading to improvements in downstream performance across multiple domains [11].
-With the rise of cross-modal contrastive models, such as CLIP [44], which align images and text in a shared embedding space, new challenges in latent space geometry have emerged.
-A notable issue is the modality gap [35], where embeddings from different modalities, such as images and text, are separated in the shared latent space.
-Moreover, the narrow cone effect was observed [35, 47], where features occupy only a limited portion of the angular space.
+image classification (Chen et al., 2020; He et al., 2020), natural language processing (Gao et al., 2021; Kim et al., 2021), 3D analysis (Afham et al., 2022; Xie et al., 2020) and more.
+The latent space induced by contrastive learning has been widely explored (Arora et al., 2019; Ji et al., 2023; Wang et al., 2022; Wang & Isola, 2020), often conceptualized as a normalized hypersphere (Wang & Isola, 2020; Liang et al., 2022).
+Alignment and uniformity (Wang & Isola, 2020) are key properties of the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss (Chen et al., 2020).
+Optimizing alignment and uniformity was shown to be crucial for preserving rich semantic structures in the latent space, leading to improvements in downstream performance across multiple domains (Fahim et al., 2024).
+With the rise of cross-modal contrastive models, such as CLIP (Radford et al., 2021), which align images and text in a shared embedding space, new challenges in latent space geometry have emerged.
+A notable issue is the modality gap (Liang et al., 2022), where embeddings from different modalities, such as images and text, are separated in the shared latent space.
+Moreover, the narrow cone effect was observed (Liang et al., 2022; Schrodi et al., 2024), where features occupy only a limited portion of the angular space.
 One of the main challenges in multimodal contrastive learning is of obtaining high-quality pairs.
-Web-scale datasets may include mismatched positive pairs [9, 12, 38, 50] or mislabeled negative pairs that are actually positive, referred to as false negatives [3, 33, 56].
-Numerous approaches have emerged to address this challenge, such as by identifying and introducing hard negative examples [4, 8, 46, 25].
+Web-scale datasets may include mismatched positive pairs (Chun et al., 2022; Gadre et al., 2024; Maini et al., 2023; Wang et al., 2023) or mislabeled negative pairs that are actually positive, referred to as false negatives (Byun et al., 2022; Li et al., 2022; Yang et al., 2022).
+Numerous approaches have emerged to address this challenge, such as by identifying and introducing hard negative examples (Byun et al., 2024; Chuang et al., 2020; Robinson et al., 2020; Kalantidis et al., 2020).
 Our observations are that false negatives appear to play a significant role in forming the geometry of CLIP’s latent space.
@@
 Let $\mathcal{X}_{i}\subset\mathcal{X}$ be the image subspace and $\mathcal{X}_{t}\subset\mathcal{X}$ be the text (captions) subspace.
-We will see that they are different and in fact linearly separable.
+We will reaffirm that they are different and in fact linearly separable (Schrodi et al., 2024; Liang et al., 2022).
 Let $v\in\mathcal{X}$ be a vector in this space.
@@
 That is, for images, $\tilde{v}_{i}=v_{i}-m_{i}\,:\,v_{i}\in\mathcal{X}_{i}$ and for text $\tilde{v}_{t}=v_{t}-m_{t}\,:\,v_{t}\in\mathcal{X}_{t}$.
-Our statistical analysis and many experimental results are based on MS-COCO [36] validation set, a common standard image-text dataset.
+Our statistical analysis and many experimental results are based on MS-COCO (Lin et al., 2014) validation set, a common standard image-text dataset.
 3.2 High dimensional geometry of random vectors
@@
 3.2.1 Thin shell theory
-There is an intensive research related to the thin shell phenomenon [26, 43, 32, 23, 31].
-Definitions of log concave distributions and isotropic random vectors appear in the supplementary.
+There is an intensive research related to the thin shell phenomenon (Kannan et al., 1995; Paouris, 2006; Klartag & Lehec, 2022; Jambulapati et al., 2022; Klartag, 2023).
+Definitions of log concave distributions and isotropic random vectors appear in the Appendix.
 Since isotropic random vectors have a unit second moment for any $x(k),\,k=1,..n$, we get that the expected value of the squared Euclidean norm is
 $\operatorname{\mathbb{E}}[\|x\|^{2}]=\operatorname{\mathbb{E}}[\sum_{k=1}^{n}x(k)^{2}]=\sum_{k=1}^{n}\operatorname{\mathbb{E}}[x(k)^{2}]=n.$ (1)
-As shown for example in [43], $\operatorname{\mathbb{E}}[\|x\|^{2}]\approx\operatorname{\mathbb{E}}^{2}[\|x\|]$, the expected norm of $x$ can be approximated by
+As shown for example in (Paouris, 2006), $\operatorname{\mathbb{E}}[\|x\|^{2}]\approx\operatorname{\mathbb{E}}^{2}[\|x\|]$, the expected norm of $x$ can be approximated by
 $\operatorname{\mathbb{E}}[\|x\|]\approx\sqrt{n}.$ (2)
@@
 where the supremum is over isotropic, log-concave random vectors in $R^{n}$. Then $\sigma_{n}\leq c(\log n)^{\alpha}$, where $c$ is a universal constant.
-Recent studies have shown this bound for $\alpha=4$ [32], $\alpha=2.23$ [23] and most recently for $\alpha=\frac{1}{2}$ [31].
+Recent studies have shown this bound for $\alpha=4$ (Klartag & Lehec, 2022), $\alpha=2.23$ (Jambulapati et al., 2022) and most recently for $\alpha=\frac{1}{2}$ (Klartag, 2023).
 See more details in the above papers and the references therein.
@@
 This part is completely data-driven without any prior assumptions related to the training process.
-We focus on the raw CLIP embedding, which is the output of the encoder before $L_{2}$ normalization, i.e.
+We focus on the primary CLIP embedding, which is the output of the encoder before $L_{2}$ normalization, i.e.
 before projection onto the unit hypersphere.
@@
 It basically “flattens” the original geometry artificially, in a manner which is hard to analyze.
-More details and statistical data are provided in the Supp.
-Let us first examine the known modality gap phenomenon [35] in the raw embedding.
+More details and statistical data are provided in the Appendix.
+Figure 3: Separability of features (left) and 10 most significant features $\ell$ for image and text, with high absolute mean, compared to the feature’s standard deviation.
+Figure 4: Statistics of image and text features after mean subtraction.
+Top: The first 10 features for image (top) and text (bottom).
+Bottom: Histograms of $\|\tilde{v}\|$ for images and text, showing a thin-shell phenomenon with no volume below a threshold, typical for high dimensions.
+Let us first examine the known modality gap phenomenon (Liang et al., 2022) in the primary embedding.
 In Fig.
@@
 For feature 93, for instance, the KL-divergence between the distributions is $\approx 301$ (a value above 1 implies a considerable deviation between the distributions).
-It was previously shown in [48, 11] that image and text can be separated linearly.
+It was previously shown in (Shi et al., 2023; Fahim et al., 2024; Schrodi et al., 2024) that image and text can be separated linearly.
 We find there are actually 9 features which serve as sort of “tags” for image and text.
@@
 $Sep(\ell)=\frac{|m_{i}(\ell)-m_{t}(\ell)|}{\sqrt{\operatorname{var}(v_{i}(\ell))+\operatorname{var}(v_{t}(\ell))}}.$ (5)
-Figure 3: Separability of features (left) and 10 most significant features $\ell$ for image and text, with high absolute mean, compared to the feature’s standard deviation.
 A plot of the features sorted by separability is shown in Fig.
 3 (left).
-In Fig.
-2 (right) it is shown that the modalities are linearly separable (with $100\%$ accuracy) using only two such tag features (93 and 134), based on a linear SVM classifier (decision boundary shown in blue).
-We can thus state the following property (which holds exactly at least for MS-COCO):
+Fig.
+2 (right) shows that the modalities are linearly separable (with $100\%$ accuracy) using only two such tag features (93 and 134), based on a linear SVM classifier (decision boundary shown in blue).
+We can thus state the following property (which holds exactly for MS-COCO):
 {mdframed}
 [ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 1: Image and text reside on separate subspaces, $\mathcal{X}_{i}\cap\mathcal{X}_{t}\approx\emptyset$.
-Figure 4: Illustration of statistics of image and text features after subtracting the mean for each modality.
-Top, histograms of 10 first features of image (first row) and text (second row).
-Bottom, histograms of $\|\tilde{v}\|$ of images and of text.
-Although each feature separately has most of the volume near zero, the distribution of $\|\tilde{v}\|$ has no volume below some threshold, indicating a thin shell phenomenon, typical for high dimensions.
 In Fig.
 4, we show some statistics of the features of $\tilde{v}_{i}$ and $\tilde{v}_{t}$ (where the mean is subtracted).
-The first 10 features in each vector are shown for both modalities.
+To get impression, the first 10 features in each vector are shown for both modalities.
 The distribution appears smooth, unimodal, with peak around zero.
 The norm $\|\tilde{v}\|$, however, is distributed within a small range (thin shell) such that there is no mass near zero.
+Figure 5: Normalized histograms of feature variance (left) show a long tail, indicating an ellipsoid rather than a hypersphere.
+Off-diagonal dominance (Eq.
+6) suggests strong feature correlations, implying a tilted ellipsoid.
 We can further check the validity of Eq.
@@
 We can therefore conclude: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 2: The mass of each modality is concentrated within a thin shell, with zero mass near the mean of the distribution.
-Figure 5: Normalized histograms of the variance of features (two left plots) indicates a long tail of high variance, hence an ellipsoid rather than an hypersphere.
-Off diagonal dominance, Eq.
-6, implies certain strong feature correlations, or a tilted ellipsoid.
 Let us now investigate the geometry of each shell.
@@
 Thus we can state: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 5: The ellipsoids are not centered near the origin.
-A linear transformation can be applied to the embedding to transform the ellipsoid into a sphere.
-This process is called whitening [10, 28].
-Analysis of whitened CLIP is in the supplementary, providing indications that multivariate normal distribution is a rather crud approximation.
 5 Loss behavior on a double-ellipsoid
-In this section, we validate that a shifted double-ellipsoid structure achieves optimality in terms of the CLIP contrastive learning loss.
+In this section, we validate that a non-origin-centered double-ellipsoid structure achieves optimality in terms of the CLIP contrastive learning loss.
 For a batch containing $M$ image-text pairs, we denote by $\bar{v}_{i}^{j}=\frac{v^{j}_{i}}{\|v^{j}_{i}\|}$ and $\bar{v}_{t}^{j}=\frac{v^{j}_{t}}{\|v^{j}_{t}\|}$ the normalized image and text features of the $j$-th pair in the batch respectively.
-The multi-modal learning loss used in CLIP is the normalized temperature-scaled cross entropy loss (NT-Xent), a variation of InfoNCE [42] loss:
+The multi-modal learning loss used in CLIP is the normalized temperature-scaled cross entropy loss (NT-Xent), a variation of InfoNCE (Oord et al., 2018) loss:
 $\ell_{clip}:=-\frac{1}{2}\mathop{\mathbb{E}}_{\begin{subarray}{c}j,k\in M\end{subarray}}\left[\log\frac{e^{{\bar{v}_{t}^{j\top}}\bar{v}^{j}_{i}/\tau}}{\sum_{j}e^{\bar{v}_{t}^{j\top}\bar{v}_{i}^{k}/\tau}}+\log\frac{e^{{\bar{v}_{t}^{j\top}}\bar{v}^{j}_{i}/\tau}}{\sum_{j}e^{\bar{v}_{t}^{k\top}\bar{v}_{i}^{j}/\tau}}\right].$ (7)
-As observed by [51], the loss can be decomposed into two terms: (1) Alignment, which encourages high cosine similarity for positive pairs, and (2) Uniformity, which encourages low cosine similarity among negative samples.
+As observed by (Wang & Isola, 2020), the loss can be decomposed into two terms: (1) Alignment, which encourages high cosine similarity for positive pairs, and (2) Uniformity, encourages low cosine similarity among negative ones.
 $\begin{split}\ell_{clip}&:=-\overbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}j\in M\end{subarray}}[\bar{v}_{t}^{j\top}\bar{v}^{j}_{i}/\tau]}^{\text{alignment}}+\\ &\overbrace{\mathop{\mathbb{E}}_{\begin{subarray}{c}k\in M\end{subarray}}\left[\frac{1}{2}\log\sum_{j=1}^{M}e^{\bar{v}_{t}^{j\top}\bar{v}^{k}_{i}/\tau}+\frac{1}{2}\log\sum_{j=1}^{M}e^{\bar{v}_{t}^{k\top}\bar{v}_{i}^{j}/\tau}\right]}^{\text{uniformity}}.\end{split}$ (8)
+Figure 6: Loss vs.
+embedding center position.
+The parameter $\alpha$ controls the embedding center (Eq.
+9, with $\alpha=0$ as the current non-origin-centered CLIP position).
+(Top).
+The unified loss balances uniformity and alignment optimally for non-origin-centered positions.
+(Bottom).
+The loss increases for misclassified instances and decreases for well-classified ones, with balanced accuracy at $\alpha\approx 0$.
 To empirically analyze the uniformity and alignment terms in Eq.
@@
 The overall loss is further separated into correctly classified, misclassified, and combined cases; the union of the correct and misclassified is equivalent to both, and they are mutually exclusive.
-Figure 6: Loss with respect to position of center of embedding.
-The parameter $\alpha$ controls the center of embedding, see Eq.
-9 ($\alpha=0$ is the current non-origin-centered position of CLIP).
-(Top).
-For non-origin-centered position, the unified loss is optimally balancing between uniformity and alignment.
-(Bottom).
-The loss is ascending when only misclassified instances are examined, and descending for well-classified.
-The overall accuracy, contains both, well balanced at $\alpha\approx 0$.
 In this experiment, we examine different values of the mean value of the image embedding.
@@
 $v_{i}^{j^{\prime}}=v_{i}^{j}-\alpha\cdot m_{i}\quad\forall j\in M.$ (9)
-The values of $v_{t}$ remain unchanged.
+The values of $v_{t}$ remain unchanged. Unlike the Embedding Shift Experiment in (Liang et al., 2022), here, the modalities are not directly shifted to each other, but to the origin.
 The results show that the loss for correctly classified samples decreases monotonically with the shift toward the origin (i.e.
-that for perfect alignment as assumed for example by [35], shifting to the origin would be preferable).
+that for perfect alignment as assumed for example by (Liang et al., 2022), shifting to the origin would be preferable).
 Conversely, the loss for misclassified samples increases.
@@
 This aligns with the current CLIP embedding, though some deviation is expected, as the MS-COCO validation set is only an approximation of the full training set.
-For completeness, the supplementary material includes the same experiment with the text ellipsoid shifted instead of the image, showing consistent behavior. To conclude:
-{mdframed}
-[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 6: CLIP’s contrastive loss is optimized for shifted ellipsoids, balancing alignment and uniformity for both correctly classified and misclassified instances.
+For completeness, the Appendix includes the same experiment with the text ellipsoid shifted instead of the image, showing consistent behavior.
+To conclude: {mdframed}[ middlelinecolor = black, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Property 6: CLIP’s loss is optimized for non-origin-centered ellipsoids, balancing alignment and uniformity for both correct and misclassified instances.
+Figure 7: Top: Example of segmentation score blur (right), common in semantic segmentation, as object-membership uncertainty increases.
+Bottom: Similarity histograms of normally distributed samples for the mean vector (blue) and the furthest vector from the mean (”extreme”, orange).
+Results are shown for a sphere centered near the origin (left) and one centered far from the origin (right).
+In contrastive learning, blur can be controlled by adjusting the sphere’s offset.
+Embedding vectors closer to the center induces blur, while positioning them farther away sharpens the response.
 6 False negatives and conformity
-We show here how the geometric structure of the embedding, found in previous sections, is advantageous when false negatives exist.
-We then propose the notion of conformity, which appears to play a major role in forming the distribution of the latent space.
-In contrastive learning, a well known problem is the phenomenon of false negatives.
-Essentially, concepts which are not dedicated pairs have similar meaning.
-Thus, they do not represent well negative examples and should not be embedded far apart.
-This problem appears for both single and multi-modality cases.
-There have been several suggestions to solve this problem by introducing new learning procedures or proposing alternative contrastive losses [4, 8].
-In the case of CLIP, however, the training is performed with a contrastive loss which does not take into consideration directly the false negative problem.
-We claim that the issue is mitigated, to some extent, by a specialized embedding.
-In classification and segmentation, as uncertainty grows, the network yields softer values, representing the lower membership probability to the class.
-In Fig.
-7 (top) a typical segmentation score of semantic segmentation is shown, where the boundary of the sheep is blurred, due to reduced confidence.
-We refer to this as semantic blur.
-For a contrastive network, when false negatives are suspected we expect lower confidence and a more blurred response.
-In the case of a high dimensional sphere centered at the origin, blurring is not trivial.
-Small perturbations yield large disparities in cosine distance.
-We show below how a sphere not centered at the origin can mitigate this.
-Figure 7: Top: blur of the segmentation score (right), typical in sematic segmentation, as the uncertainty of object-membership grows.
-Bottom: the similarity histograms of normally distributed samples are shown for the mean vector (blue) and a vector furthest from the mean (“extreme”, orange).
-The plots are done for a sphere centered near the origin (left) and for one with a center far from the origin (right).
-See full details of the experiment in the text.
-In the case of contrastive learning, blur can be achieved, as uncertainty grows, by having a sphere centered far from the origin.
-The desired blur can be controlled by positioning the embedded vector closer to the sphere center (blur) or farther from it (sharp).
-[…]
+We demonstrate how the embedding geometry discussed earlier provides advantages in handling false negatives.
+Additionally, we introduce the concept of conformity, which plays a major role in forming the latent space distribution.
+A well-known challenge in contrastive learning is the presence of false negatives—pairs with similar meanings that are not dedicated pairs.
+Such samples should not be embedded far apart, as they fail to represent true negatives effectively.
+This issue arises in both single- and multi-modality settings and has been addressed by proposing new training procedures or alternative contrastive losses (Byun et al., 2024; Chuang et al., 2020).
+In CLIP, training uses a contrastive loss that does not explicitly address false negatives.
+However, we argue that this issue is partially mitigated by the embedding geometry.
+In classification and segmentation tasks, uncertainty typically results in softer predictions that reflect lower class membership probabilities.
+For example, Fig.
+7 (top) illustrates a segmentation score where reduced confidence blurs the sheep’s boundary, a phenomenon we term semantic blur.
+Figure 8: High and low conformity of MS-COCO.
+Low-conformity images often depict unique, distinguishable individuals or objects, whereas high-conformity images capture common scenes that could be found anywhere.
+For contrastive networks, when false negatives are present, we expect lower confidence and a blurred response.
+On a high-dimensional sphere centered at the origin, such blurring is challenging, as small perturbations lead to large changes in cosine distance.
+We show that shifting the sphere away from the origin can effectively mitigate this issue.
+Concurrently, and closely related, Schrodi et al.
+(2024) discuss the relationship between entropy and the modality gap.
+Blur through a non-origin centered sphere.
+Fig.
+7 (bottom) illustrates the difference between origin-centered and non-origin-centered spheres through an experiment.
+We draw 1000 random vectors $v^{j}\in\mathbb{R}^{512}$, where each element follows an independent Gaussian distribution with unit standard deviation.
+In the first experiment (Fig.
+7, bottom left), the sphere is centered at the origin with an empirical mean $m$ close to zero.
+The blue histogram shows $\cos(m,v^{j})$ for $j=1,\dots,1000$.
+We then identify the furthest vector from $m$, $v^{\text{far}}=\arg\min\cos(m,v^{j})$, and plot the histogram of $\cos(v^{\text{far}},v^{j})$ (orange), excluding $v^{\text{far}}$.
+In the second experiment (Fig.
+7, bottom right), the sphere is centered at $(10,5,5,0,0,\dots)$, modeling three dominant features with a mean distinctly far from zero.
+The same trial is repeated.
+The results highlight a significant difference: for an origin-centered sphere, the distributions of cosine similarity for the mean and the extreme vector are similar.
+In contrast, for a non-origin-centered sphere, the mean vector exhibits much higher average similarity.
+This allows the network to embed vectors with uncertainty closer to the mean, enabling semantic blur—reduced contrast in the response.
+This analysis, supporting a non-zero mean, leads to the following prediction:
 {mdframed}
 [ middlelinecolor = blue, middlelinewidth = 0.5pt, backgroundcolor = gray!20, roundcorner = 4pt, ] Prediction 1: Common themes, which occur more frequently in the training set, are expected to be embedded in closer proximity to the mean vector.
-Figure 8: High and low conformity of natural images from MS-COCO [36].
-Low-conformity images often depict unique, distinguishable individuals or objects, whereas high-conformity images capture common scenes that could be found anywhere.
+Figure 9: Conformity.
+Estimated conformity $\hat{C}$, Eq.
+11, against conformity $C$, Eq.
+10, on MS-COCO (Lin et al., 2014).
+The correlation is almost perfect.
+We can thus use the proposed estimated conformity reliably to quantify how common a sample is.
+More exotic captions have lower conformity (all examples are of eight words).
+Figure 10: Conformity Differences. The conformity distributions of text and image modalities differ, as a common image may be described by a unique caption, and vice versa.
+Figure 11: Modality Gap matches conformity distributions.
+The parameter $\alpha$ controls the embedding offset from the origin (as shown in Fig.
+6).
+When $\alpha\approx 0$, i.e., the trained setting, image and text conformity distributions align well, with $KL_{\alpha=0}\approx 0.14$ indicating good distribution matching.
 6.1 Conformity
-In order to validate our prediction, we first formalize the term common themes, by defining a new notion, termed conformity.
+To validate Prediction 1, we first formalize the term common themes, by defining a new notion, termed conformity.
 Definition 1 (Conformity).
@@
 To provide more intuition, we present examples of high and low conformity from MS-COCO in Fig.
-8, as well as on ImageNet-a [21] and ImageNet-R [20] in the Supp.
+8, as well as on ImageNet-a (Hendrycks et al., 2021b) and ImageNet-R (Hendrycks et al., 2021a) in the Appendix.
 Following our prediction above, we propose a surrogate measure of conformity (which is much faster to compute).
@@
 $\hat{C}(v^{j})=a\cdot\cos(m,v^{j})+b,$ (11)
-where $a$ and $b$ are scalars determined by the specific embedding.
-Figure 9: Conformity.
-Estimated conformity $\hat{C}$, Eq.
-11, against conformity $C$, Eq.
-10, on MS-COCO [36].
-The correlation is almost perfect.
-We can thus use the proposed estimated conformity reliably to quantify how common a sample is.
-Employing $\hat{C}$ is much faster, using a single mean vector $m$, without the need to store an entire set of representative image-text pairs.
-We also show a few caption examples of high and low conformity, with the same number of words (8).
-More exotic captions have lower conformity.
+where $a$ and $b$ are scalars determined by the embedding.
 In Fig.
@@
 6.2 Modality gap assists in distribution matching
-We now aim to provide a reason that can justify the presence of the well known modality gap [35]. Our rationale for that phenomenon is as follows.
-Figure 10: Conformity Differences. The conformity distributions of text and image modalities differ, as a common image may be described by a unique caption, and vice versa.
+We now aim to provide a reason that can justify the presence of the well known modality gap (Liang et al., 2022).
+Our rationale for that phenomenon is as follows.
 The same incentive of having a mean not centered at the origin applies for both image and text modalities.
@@
 6.
-We show that the best distribution match is near $\alpha=0$, that is, in the current embedding of CLIP.
-Figure 11: Matching the distribution of conformity for image and text by introducing a modality gap.
-The parameter $\alpha$ controls the shift from the origin for images (as in Fig.
-6).
-Both image and text conformity distributions are aligned when $\alpha\approx 0$, i.e.
-without any shift from the trained setting.
-$KL_{\alpha=0}\approx 0.14$ yields a good distribution matching.
-Figure 12: Vertical SLERP (vSLERP) - Optimization-free, spontaneous semantic editing.
-Images interpolated with vSLERP preserve the same object with slight pose variations, while roughly maintaining backgrounds.
-The interpolation magnitude is controlled by $\alpha$.
+We show that the best distribution match is near $\alpha=0$, i.e., in the current embedding of CLIP.
+Figure 12: Vertical SLERP (vSLERP) enables optimization-free, semantic editing.
+Interpolated images preserve the object with pose variations and roughly maintain backgrounds, with interpolation magnitude controlled by $\alpha$.
 Figure 13: Conformity analysis of captioning and image synthesis.
-Image Synthesis (top): Glide [41] tends to generate more common images, often struggling with fine-grained details, whereas unCLIP [45] generates images with more delicate details, aligning more closely with the distribution of natural images.
-Though there is still room for improvement.
-Captioning (bottom): ClipCap [40] generates captions with higher conformity (i.e., more common descriptions) compared to human-annotated captions.
-In contrast, Caption Reward [7] produces captions with lower conformity, even surpassing human annotations in terms of uniqueness.
+Image Synthesis (top): Glide generates more common images with less fine detail, while unCLIP creates more detailed images closer to natural distributions.
+Captioning (bottom): ClipCap produces more common captions, while Caption Reward generates more unique captions, even surpassing human annotations.
 7 Applications
 7.1 Conformity as a measure of expressiveness
-We propose to use (estimated) conformity as a metric to assess the richness of generative methods.
-Specifically, we measure the conformity of images generated from MS-COCO captions by unCLIP [45] and Glide [41], visualizing the results in Fig.
+We propose using conformity as a metric to assess generative method diversity.
+We measure conformity in images generated from MS-COCO captions by unCLIP (Ramesh et al., 2022) and Glide (Nichol et al., 2021), as shown in Fig.
 13.
-Glide-generated images exhibit high conformity, indicating less detail and diversity, whereas unCLIP images are more varied and detailed.
-However, both models fall short of matching the diversity found in real images.
-Similarly, we evaluate captioning methods by measuring conformity in captions generated using ClipCap [40] and Caption Reward [7].
-ClipCap tends to produce common and trivial captions, while Caption Reward generates captions with diversity surpassing even human annotations.
+Glide-generated images exhibit high conformity, indicating low detail and diversity, while unCLIP images are more varied and detailed.
+Both models, however, lack the diversity seen in real images.
+Similarly, we evaluate captioning methods by measuring conformity in captions generated by ClipCap (Mokady et al., 2021) and Caption Reward (Cho et al., 2022).
+ClipCap produces common captions, while Caption Reward generates diverse captions that even surpass human annotations.
 7.2 Unguided, training-free semantic generation
-The unCLIP framework [45] introduces an image interpolation technique using spherical linear interpolation (SLERP) to transform a source image into a target image gradually.
-While this method produces visually appealing results, it often fails to preserve the same instance throughout interpolation, instead generating random instances.
-In Fig. 12, we show images generated by an extension of SLERP, which we term as vertical SLERP (vSLERP), defined as:
+The unCLIP framework (Ramesh et al., 2022) introduces an image interpolation technique using spherical linear interpolation (SLERP) to transform a source image into a target image gradually.
+While this method produces visually appealing results, it often fails to preserve the same instance along interpolation, instead generating random instances.
+In Fig. 12, we show images generated by an extension of SLERP, which we term as vertical SLERP (vSLERP):
 $\textit{vSLERP}(v^{j},v^{k},\theta_{0},\alpha)=\textit{SLERP}(v^{j}-\alpha m,v^{k}-\alpha m,\theta_{0})+\alpha m$ (12)
 For brevity, $m_{i}$ and $v_{i}$ are referred to as $m$ and $v$.
-With a fixed $\Theta=\Theta_{0}$, adjusting $\alpha$ allows controlled manipulation of the same instance (values in the range $-2\leq\alpha\leq 2$ tend to produce visually appealing images).
-This approach parallels real-image editing techniques; however, unlike methods relying on text inversion [16, 13, 39] or test-time optimization [27], which are computationally heavy, vSLERP requires no training or optimization, thus, highly efficient.
+With a fixed $\Theta=\Theta_{0}$, adjusting $\alpha$ allows controlled manipulation of the same instance.
+This approach parallels real-image editing techniques; however, unlike methods relying on text inversion (Han et al., 2024; Gal et al., 2022; Mokady et al., 2023) or test-time optimization (Kawar et al., 2023), which are computationally heavy, vSLERP requires no training or optimization, thus, highly efficient.
 8 Discussion and Conclusion
-This paper analyzes the raw CLIP embeddings before projection onto a unit hypersphere, revealing that each modality’s distribution forms a distinct, shifted ellipsoid.
-These ellipsoids differ in properties such as their center of mass and radii and are disjoint, hence the modality gap phenomenon [35].
-Upon projection onto a unit hypersphere, they occupy only a limited angular space, which appears as a narrow cone effect [35, 47].
-We introduce the concept of conformity, a measure of similarity with the entire set, to gain further insights into the double-ellipsoid geometry.
-We observe that each modality has a unique conformity distribution, with optimal matching achieved when the ellipsoids are positioned away from the center.
-Using this framework, we evaluate the diversity of captioning and image synthesis methods and propose a training-free interpolation technique, vertical SLERP (vSLERP), for specific object interpolation.
+The paper examines the primary CLIP embedding, prior to projection onto the unit sphere, revealing that each modality forms a distinct, shifted ellipsoid with unique centers and radii.
+This geometry is the source of the modality gap and narrow cone phenomena (Liang et al., 2022; Schrodi et al., 2024; Afham et al., 2022), previously observed on the unit sphere embedding.
+We introduced conformity, a measure of similarity of an instance with an entire representative set.
+Our analysis shows that each modality exhibits a unique conformity distribution, with optimal alignment achieved when the ellipsoids are shifted from the origin.
+This provides a useful tool for assessing the diversity of captioning and image synthesis methods.
+Finally, we propose vertical SLERP (vSLERP), a training-free interpolation technique for specific object interpolation.
+Acknowledgements
+We would like to acknowledge support by the Israel Science Foundation (Grant 1472/23) and by the Ministry of Science and Technology (Grant No. 5074/22).
+Impact Statement
+Our work advances machine learning by improving the geometrical understanding of CLIP’s latent space.
+The findings may influence downstream tasks, though specific societal consequences do not need further emphasis here.
+Appendix A Enlraged Visualizations
+In Figure 15 and Figure 16, we provide the same visualizations as in the main paper, but enlraged, to enhance visibility.
+CLIP of higher dimension.
+We also show some results for CLIP with ViT-L/14 encoders, $n=768$.
+In Figure 17 we show the distinct different statistics of image and text, mostly appearing in several pronounced features.
+Here as well, linear separation (100% classification accuracy) can be reached with only two features.
+In Figure 18 we show that the embedding can also be modeled as two separate thin shell ellipsoids for image and text.
+Appendix B Statistical Analysis
+We provide here the definitions of log concave distributions and isotropic random vectors, notions which are used in Section 4 of the main paper.
+Definition 3 (Log concave distribution).
+A log concave distribution in $R^{n}$ has a density $p$ which admits, $\forall x,y\in R^{n},\lambda\in[0,1]$,
+$p(\lambda x+(1-\lambda)y)\geq p(x)^{\lambda}p(y)^{1-\lambda}.$
+The above definition is equivalent to stating that the logarithm of the density function is concave $\log p(\lambda x+(1-\lambda)y)\geq\lambda\log p(x)+(1-\lambda)\log p(y)$.
+Many well-known distributions admit this property, such as normal and multivariate normal distributions, exponential, Laplace, chi, Dirichlet, gamma and more.
+Definition 4 (Isotropic random vector).
+A random vector $x\in R^{n}$ is isotropic if $\operatorname{\mathbb{E}}[x]=0$ and $\Sigma=I$, where $\Sigma$ is the covariance matrix of $x$ and $I$ is the identity matrix.
+Figure 14: Norm distribution. While norm magnitudes are disregarded during training due to the normalization inherent in cosine similarity, they still capture meaningful semantic information.
+We give below additional analysis related to applying a linear transformation that turns each ellipsoid into a sphere.
+This process is termed sphering or whitening.
+For lack of space, this part did not get into the main paper.
+However, we believe this analysis is of sufficient merit to be presented here.
+Figure 15: Enlarged plots from Section 4.
+Figure 16: Enlarged plots from Section 4.
+Figure 17: Enlarged plots for CLIP embedding of $n=768$.
+There are dominant features with clearly different distribution between image and text.
+Both modalities can be separated (with perfect accuracy) by a linear SVM classifier based on only 2 features.
+With respect to separability (bottom), there are 20 features with value above 1.
+Figure 18: CLIP $n=768$, thin shell phenomenon. We can observe similar geometry (as in the case of $n=512$) of two tilted ellipsoids, one for each modality, not centered at the origin.
+Appendix C Additional Experiments and Visualizations
+C.1 Conformity
+High- and Low-Conformity Images.
+We provide additional visualizations of high- and low-conformity images across various datasets.
+Figure 19 illustrates examples of sketches from ImageNet-R, while Figure 20 showcases examples from ImageNet-A.
+Both datasets contain out-of-distribution examples: ImageNet-A emphasizes natural adversarial images, while ImageNet-R features renditions of objects, such as origami or sketches.
+From these visualizations, we observe that high-conformity images tend to contain less information.
+Sketches are simpler, and natural images often feature large uniform backgrounds or repetitive structures.
+In contrast, low-conformity images frequently include substantial text, while natural images exhibit collages of objects with unique or diverse colors.
+C.2 Reaffirming loss and conformity matching experiments
+We revisit the loss experiment presented in Fig.
+6 of the main paper and the conformity matching experiment shown in Fig.
+11.
+To further validate our findings, we conduct these experiments under two alternative settings.
+First, we shift the text ellipsoid instead of the image ellipsoid, applying the following transformation:
+$v_{t}^{j^{\prime}}=v_{t}^{j}-\alpha\cdot m_{t}\quad\forall j\in M,$ (13)
+where the values of $v_{i}$ remain unchanged. The results of this experiment are presented in Figure 21.
+In the second setting, we align both the image and text ellipsoids at the origin by applying the following transformations:
+$v_{t}^{j^{\prime}}=v_{t}^{j}-\alpha\cdot m_{t},\quad v_{i}^{j^{\prime}}=v_{i}^{j}-\alpha\cdot m_{i}\quad\forall j\in M.$ (14)
+Here, for $\alpha=0$, the ellipsoids remain in their optimal positions after training, while for $\alpha=1$, both ellipsoids are shifted to the origin as in Figure 22.
+Both experiments reaffirm that the current positioning of the ellipsoids yields optimal results in terms of loss and conformity matching.
+These findings further support our claims across different alignment scenarios.
+C.3 vSLERP
+Here, we provide additional examples of vSLERP, shown in Figure 23 and Figure 24.
+As discussed in the main paper, the standard SLERP process typically generates interpolated images representing different objects or individuals.
+In contrast, our proposed vSLERP method produces diverse variations of the same object.
+Figure 19: High and low conformity of sketches from ImageNet-R.
+Images with high conformity tend to be simpler and cleaner, while low-conformity images often feature complex details covered by large portions of text descriptions.
+Figure 20: Conformity on ImageNet-a.
+It is possible that high conformity images are with more unique colors, perhaps contains people or text, whereas low conformity images tends to contain low amount of information.
+Figure 21: Shifting text ellipsoid only. Conformity distribution matching and loss experiments when shifting text ellipsoid only as in Equation 13
+Figure 22: Shifting both ellipsoids. Conformity distribution matching and loss experiments when shifting both text and image ellipsoids as in Equation 14.
+Figure 23: vSLERP lamp to vase.
+Figure 24: vSLERP Kevin Durant to Lebron James.

### v3

@@
 A new notion of conformity is introduced, which measures the average cosine similarity of an instance to any other instance within a representative data set.
-We show this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector.
+We prove this measure can be accurately estimated by simply computing the cosine similarity to the modality mean vector.
 Furthermore, we find that CLIP’s modality gap optimizes the matching of the conformity distributions of image and text.
@@
 Leveraging this deeper understanding, we introduce a new definition of concept conformity, quantifying how close a sample resides with respect to all others.
-We find a remarkable correlation between conformity and cosine similarity to the mean vector (Pearson correlation: 0.9998 for MS-COCO).
+We prove that there is a proportion between conformity and cosine similarity to the mean vector (See proof in Supp.
+C1, and empirically with Pearson correlation: 0.9998 for MS-COCO).
 Furthermore, we show that the distribution of conformity differs between modalities, with CLIP’s ellipsoid alignment offering a plausible explanation for the modality gap.
@@
 11, against conformity $C$, Eq.
-10, on MS-COCO (Lin et al., 2014).
+13, on MS-COCO (Lin et al., 2014).
 The correlation is almost perfect.
@@
 where $a$ and $b$ are scalars determined by the embedding.
-In Fig.
+In Appendix C1 we prove this correlation under the thin-shell assumption, and in Fig.
 9, $C$ versus $\hat{C}$ are plotted for the entire MS-COCO set, for both image and text embeddings.
-A close to perfect correlation is obtained, with Pearson correlation of 0.9998 for both image and text where $a=1.411$, $b=-0.008$ for text and $a=1.461$, $b=-0.002$ for images.
+A close to perfect correlation is obtained, with Pearson correlation of 0.9998 for both image and text where $a=1.411$, $b=-0.008$ for text and $a=1.461$, $b=-0.002$ for images, validating with close to perfect alignment with the rigorous mathematic derivation.
 6.2 Modality gap assists in distribution matching
@@
 However, we believe this analysis is of sufficient merit to be presented here.
+Appendix C Additional Experiments and Visualizations
+C.1 Close relations between conformity and surrogate-conformity
+We show below the validity of our conformity approximation under the thin-shell assumption.
+Proposition 1.
+Let $S=\{v^{1},\dots,v^{N}\}$ be a set of $N$ vectors in $\mathbb{R}^{F}$ exhibiting the thin-shell phenomenon, i.e.,
+$\|v^{i}-\bar{v}\|\approx R\quad\text{for all }i,$
+where $\bar{v}=\frac{1}{N}\sum_{j=1}^{N}v^{j}$ is the sample mean and we use the Euclidean norm $\|v^{i}\|^{2}=\sum_{k}(v^{i}_{k})^{2}$. Then, for any $v^{j}\in S$, the following approximation holds:
+$\mathbb{E}_{v^{j}\in S}[\cos(v^{i},v^{j})]\approx A\cdot\cos(v^{i},\bar{v}),$ (13)
+where $A\approx\frac{\sqrt{\mu_{\text{norm}}^{2}+R^{2}}}{\mu_{\text{norm}}}$, $\mu_{\text{norm}}=\|\bar{v}\|$ and the symbol $\approx$ represents the shell approximation (which becomes more accurate as the width of the shell decreases) and approximate orthogonality between a random vector and the mean vector.
+Proof.
+We start by expanding the left-hand side:
+$\mathbb{E}_{v^{j}\in S}[\cos(v^{i},v^{j})]=\frac{1}{N}\sum_{j=1}^{N}\frac{v^{i}\cdot v^{j}}{\|v^{i}\|\cdot\|v^{j}\|}.$
+Writing explicitly the inner-product we have:
+$\frac{1}{N\cdot\|v^{i}\|}\sum_{j=1}^{N}\frac{1}{\|v^{j}\|}\sum_{k=1}^{F}v^{i}_{k}v^{j}_{k}.$
+Now, consider the right-hand side of Equation (13):
+$\cos(v^{i},\bar{v})=\frac{v^{i}\cdot\bar{v}}{\|v^{i}\|\cdot\|\bar{v}\|}=\frac{1}{\|v^{i}\|\cdot\mu_{\text{norm}}}\sum_{k=1}^{F}v^{i}_{k}\left(\frac{1}{N}\sum_{j=1}^{N}v^{j}_{k}\right)=\frac{1}{N\cdot\|v^{i}\|\cdot\mu_{\text{norm}}}\sum_{j=1}^{N}\sum_{k=1}^{F}v^{i}_{k}v^{j}_{k}.$
+Observe that the only difference between the two expressions lies in the difference between $\mu_{\text{norm}}$ and $\|v^{j}\|$.
+We show below that under the thin-shell assumption $\|v^{j}\|\approx\sqrt{R^{2}+\mu_{\text{norm}}^{2}}$.
+Let us define by $z^{j}$ the difference vector between a vector $v^{j}$ and the mean vector $\bar{v}$, that is $z^{j}=v^{j}-\bar{v}$. Then,
+$\|v^{j}\|^{2}=\|z^{j}+\bar{v}\|^{2}=\|z^{j}\|^{2}+2z^{j}\cdot\bar{v}+\|\bar{v}\|^{2}.$
+In high dimensions, the inner product $z^{j}\cdot\bar{v}$ is small due to approximate orthogonality, so:
+$\|v^{j}\|^{2}\approx\|z^{j}\|^{2}+\mu_{\text{norm}}^{2}\approx R^{2}+\mu_{\text{norm}}^{2}.$
+Taking square roots:
+$\|v^{j}\|\approx\sqrt{R^{2}+\mu_{\text{norm}}^{2}}.$
+Thus, the scalar factor $A$ in Equation (13) is given by:
+$A=\frac{\mu_{\text{norm}}}{\|v^{j}\|}\approx\frac{\mu_{\text{norm}}}{\sqrt{R^{2}+\mu_{\text{norm}}^{2}}}.$
+Empirically we know for Vit-B/32 that $\mu_{norm}=7.587$ and $R\approx 7.59$, thus the mathematical derivation state that $A^{-1}=\frac{\sqrt{7.59^{2}+7.587^{2}}}{7.587}=1.414$ For images and $A^{-1}=\frac{\sqrt{5.59^{2}+5.75^{2}}}{5.75}=1.4$, very close to the empirical observations (note that the correlation is reversed in the main paper).
+C.2 Conformity
+High- and Low-Conformity Images.
+We provide additional visualizations of high- and low-conformity images across various datasets.
+Figure 19 illustrates examples of sketches from ImageNet-R, while Figure 20 showcases examples from ImageNet-A.
+Both datasets contain out-of-distribution examples: ImageNet-A emphasizes natural adversarial images, while ImageNet-R features renditions of objects, such as origami or sketches.
+From these visualizations, we observe that high-conformity images tend to contain less information.
+Sketches are simpler, and natural images often feature large uniform backgrounds or repetitive structures.
+In contrast, low-conformity images frequently include substantial text, while natural images exhibit collages of objects with unique or diverse colors.
 Figure 15: Enlarged plots from Section 4.
@@
 Figure 18: CLIP $n=768$, thin shell phenomenon. We can observe similar geometry (as in the case of $n=512$) of two tilted ellipsoids, one for each modality, not centered at the origin.
-Appendix C Additional Experiments and Visualizations
-C.1 Conformity
-High- and Low-Conformity Images.
-We provide additional visualizations of high- and low-conformity images across various datasets.
-Figure 19 illustrates examples of sketches from ImageNet-R, while Figure 20 showcases examples from ImageNet-A.
-Both datasets contain out-of-distribution examples: ImageNet-A emphasizes natural adversarial images, while ImageNet-R features renditions of objects, such as origami or sketches.
-From these visualizations, we observe that high-conformity images tend to contain less information.
-Sketches are simpler, and natural images often feature large uniform backgrounds or repetitive structures.
-In contrast, low-conformity images frequently include substantial text, while natural images exhibit collages of objects with unique or diverse colors.
-C.2 Reaffirming loss and conformity matching experiments
+C.3 Reaffirming loss and conformity matching experiments
 We revisit the loss experiment presented in Fig.
@@
 These findings further support our claims across different alignment scenarios.
-C.3 vSLERP
+C.4 vSLERP
 Here, we provide additional examples of vSLERP, shown in Figure 23 and Figure 24.